I have a couple of problems that I need help with.
1) Solve using the square root property:
5x^2 = 100
2) Solve using the square root property:
(x + 8)^2 = 81

Answer by Alan3354(69443)

have a couple of problems that I need help with.
1) Solve using the square root property:
5x^2 = 100
x^2 = 20
x = +2sqrt(5)
x = -2sqrt(5)
-------------
2) Solve using the square root property:
(x + 8)^2 = 81
x+8 = 9
x = 1
------
x+8 = -9
x = -17